OpenStudy (anonymous):

find each qoutient. write in simplest form 2 1/3 divided by 8

OpenStudy (anonymous):

2 1/3 = 7/3 and when divided by 8, it becomes 7/(3)(8) = 7/24 = 0.291666666

OpenStudy (anonymous):

\[2\frac{ 1 }{ 3 } =\frac{ 7 }{ 3 }\] \[\frac{ \frac{ 7 }{ 3 } }{ 8} = \frac{ 7 }{ 3 } * 8\] =\[\frac{ 7 }{ ? }\]

OpenStudy (anonymous):

\[\frac{ 7 }{ 24 }\]

jhonyy9 (jhonyy9):

so if 2 1/3 = (2*3 +1)/3 =(6+1)/3 =7/3 and this 7/3 you need divide by 8 so but you know that 8 is equal 8/1 so hopw you know how can you divide two fractions so hence you will get 7/3 time 1/8 = 7/(3*8) =7/24 ok ? hope so much that you have understood it right good luck bye

OpenStudy (anonymous):

2 1/3 divided by 8 change to improper fraction 7/3 divided by 8/1 get the reciprocal of 8/1 then change division to multiplication so: 7/3 times 1/8 =7/24